OpenStudy (anonymous):

Simplify -(3/4)(-8x)(4/3)(4y)(1)

OpenStudy (igreen):

Multiply the first two terms: \(-\dfrac{3}{4} \cdot -8x = 6x\) Now multiply that to the third term: \(6x \cdot \dfrac{4}{3} = 8x\) Now multiply that to the fourth term: \(8x \cdot 4y = 32xy\) Can you multiply \(32xy \cdot 1\)?

OpenStudy (igreen):

@OpiGeode

OpenStudy (anonymous):

Nope.

OpenStudy (igreen):

Well based on this rule: \(x \cdot 1 = x\) What do you think \(32xy \cdot 1\) is?

OpenStudy (anonymous):

Its okay I passed my test.

OpenStudy (igreen):

Oh..well which one did you pick? \(32xy \cdot 1 = 32xy\)